OpenStudy (anonymous):

tan^4x-sec^4x=1-2 sec^2(x) verify

OpenStudy (harsimran_hs4):

use equation generator other your question can be interpreted in many ways

OpenStudy (anonymous):

\[\tan^4x-\sec^4x=1-2 \sec^2(x)\]

OpenStudy (harsimran_hs4):

consider LHS since we need final answer in sec x we will try and substitute tanx in terms of sec x we know sec^2 x = 1 + tan^2 x tan^2 x = sec^2 x - 1 tan^4 x = (sec^2 x - 1)^2 = sec^4 x + 1 - 2sec^2 x can you proceed now??

OpenStudy (anonymous):

let me work on it and see

OpenStudy (anonymous):

LHS \[=(\sec ^{2}x-1)^{2}-\sec ^{4}x\] \[=\sec ^{4}x+1-\sec ^{4}x\] \[=1\] RHS \[=1-(\sec ^{2}x+\sec ^{2}x)\] \[=1+\sec ^{2}x-\sec ^{2}x\] \[=1\] Are these steps correct?

OpenStudy (harsimran_hs4):

no these are not correct look above at my post i have done all the steps

OpenStudy (anonymous):

im not sure how you came up with this = = sec^4 x + 1 - 2sec^2 x

OpenStudy (harsimran_hs4):

(a+b)^2 = a^2 + b^2 +2ab clear, if not think a little about it`s proof